OpenStudy (anonymous):

Help?? Simplify by combining like terms. m + 4p + 7m + 2p – 5m A 3m – 6p B 2m + 6p C 3m + 6p D 2m – 6p

OpenStudy (anonymous):

you would combine 7m+m=8m-5m=3m. Then you would combine 4p+2p=6p. That leaves you with 3m+6p which is C

OpenStudy (anonymous):

Nicely done :) (Thank you :P )

OpenStudy (anonymous):

C